OpenStudy (anonymous):

m∠HGF = 16x + 4, m∠EGF = 110°, and m∠HGE = 3x + 11. Find x.

OpenStudy (jessiegonzales):

(m<EGF) + (m<HGE) = HGF

OpenStudy (jessiegonzales):

(110)+(3X+11)=16x+4

OpenStudy (anonymous):

what if it were a different question with different letters and equations, how do i know where to put the things?

OpenStudy (jessiegonzales):

110+3x+11=16x+4

OpenStudy (jessiegonzales):

3x+121=16x+4

OpenStudy (jessiegonzales):

Now solve for X!!!

OpenStudy (anonymous):

what if it were different numbers and letters how do i know where to put each equation?

OpenStudy (jessiegonzales):

OKAY so....

OpenStudy (jessiegonzales):

THE STEPS ARE GOING TO BE THE SAME JUST DIFFERENT NUMBERS!! I WILL GIVE U AN EXAMPLE AND U PUT THEM WHERE THEY ARE SUPPOSED TO GO OK?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so basically i take the number and an equation add those together and set it equal to the equation of the adjacent angle?

OpenStudy (jessiegonzales):

yes ma'am!! can u solve for x?

OpenStudy (anonymous):

yeah let me do the math hold on

OpenStudy (jessiegonzales):

ok

OpenStudy (anonymous):

x=9!

OpenStudy (jessiegonzales):

CORRECT!!!

OpenStudy (anonymous):

oh my gosh thank you so much!!!!

OpenStudy (jessiegonzales):

NO PROBLEM!!